OpenStudy (anonymous):

\[\int\limits_{1}^{3} \frac{2t+3}{t+1}\] please integrate by definite integration. :)

OpenStudy (anonymous):

\[\int\limits_{1}^{3} \frac{2t+3}{t+1}\]

OpenStudy (anonymous):

try \(u=t+1\)

OpenStudy (anonymous):

ok. :)

OpenStudy (anonymous):

don't let the numerator throw you off, if \(u=t+1\) then \(t=u-1\) and \(2t+3)=2(u-1)+3\) etc

OpenStudy (blockcolder):

Actually, doing the division makes it simpler: \[\frac{2t+3}{t+1}=2-\frac{1}{t+1}\]

OpenStudy (anonymous):

is answer is ln(2) + 4 ?

OpenStudy (anonymous):

yes! @mrtdmr26 can you show me your solution?

OpenStudy (anonymous):

\[[ \int\limits\limits_{1}^{3}(2t)/(t+1) ] + [ \int\limits\limits_{1}^{3} 3 / (t +1) ] \] \[\int\limits\limits(2t)/(t+1) = [ -2\ln (|t+1|) + 2t - 1 + C]\] \[ \int\limits\limits 3 / (t +1) = 3\ln ( |x+1| ) + C\] when yu put the limits 1 to 3 1st Part \[4-2\ln(2)\] 2nd Part\[3\ln(2)\] \[4-2\ln(2) + 3\ln(2) = \ln(2) + 4\] That's all :)

OpenStudy (anonymous):

Thank you a thousand times @mrtdmr26 . :)